The following table shows estimates of activity times (weeks) for a project:


Activity
Optimistic
Time
Most Probable
Time
Pessimistic
Time
A 3 4.5 9
B 1.5 2 2.5
C 5 6 7
D 4 4.5 8
E 5 6 7
F 1 2.5 7
G 7 9 11

Suppose that the critical path is A-B-D-F-G. Use Appendix A to answer the questions. Do not round intermediate calculations. Round your answers to three decimal places.

  1. What is the probability that the project will be completed within 25 weeks?
Answer #1

Answer: the probability that the project will be completed within 25 weeks is 0.7208 or 72.08%

Explanation:

Activity Optimistic time-a Expected completion time-m Pessimistic time-b Expected time= (a+4*m+ b)/6 Variance, (sigma)^2= (b-a/6)^2
A 3 4.5 9 5.00 1.00
B 1.5 2 2.5 2.00 0.03
C 5 6 7 6.00 0.11
D 4 4.5 8 5.00 0.44
E 5 6 7 6.00 0.11
F 1 2.5 7 3.00 1.00
G 7 9 11 9.00 0.44

The duration of critical path ABDFG is 24 weeks,

step 1 we will find the variance of the tasks which lie on critical path 2.92
step 2 mean project time (u) of critical path is= 24.00
step 3 Required completion time is 25
step 4 standard deviation= sqrt(variance)=   1.709
step 5 because Z= (given completion time- u)/standard deviation 0.585
step 6 P(z)= using NORM.S.DIST(z,true) 0.7208

the probability that the project will be completed within 25 weeks is 0.7208

Please ask, if you have any doubts through the comment section. Do rate the answer